OpenStudy (anonymous):

Hello Guys, my best wished to you all. I just need a new account and this is my first question. Here it goes. QS: Find fundamental system and general solution of differential equation defined by .. |1/t t^2 0| |0 -2/t^2 1/t^2| |0 0 1/t | Now, as per my knowledge there are two appropriate approaches into solving this ... 1) state space method 2) solving from downward to above, first finding x3 then x2 then x1. However, I am not getting the same answer by both mehtod... MY solution is

OpenStudy (anonymous):

r = 1/t, r=1/t, r=-2/t are the eigen values for r = 1/t, first eigen vector is [1 0 0]' for r = 1/t, the second eigen vector is [1 1/t^2 3/t]' for r=-2/t, the eigen vector is [1 -3/t^3 0] oh, this is really not easy typing it all :(